OpenStudy (anonymous):

what is the equation of the line, in slope-intercept for, that goes through the point (2,-1) with m=3/8? A) y=3/8x+1/4 B) y=3/8x+7/4 C)y=3/8x-1/4 D y=3/8x-7/4

OpenStudy (mertsj):

Did you try this one? It is the same as the previous one.

OpenStudy (anonymous):

i just wanna get the right answer for sure i need help with it

OpenStudy (mertsj):

Post what you did and I will go over it for you.

OpenStudy (anonymous):

i didint do it

OpenStudy (mertsj):

Ok. I'll wait while you do it.

OpenStudy (anonymous):

u plug in m to x right?

OpenStudy (mertsj):

\[y-y _{1}=m(x-x _{1)}\]

OpenStudy (mertsj):

Better memorize that formula.

OpenStudy (anonymous):

wait so y-y1=3/8 (x-x1)? am i right so far?

OpenStudy (anonymous):

can i have the answe ive been waiting so long to have it...

OpenStudy (anonymous):

Answer is A..same methods

OpenStudy (anonymous):

(y+1)=3/8(x-2)

OpenStudy (anonymous):

okay.. thank you

OpenStudy (anonymous):

this lesson is hard for me to understand mybad